The hypotenuse of a right triangle is 9cm longer than 1 leg and 2cm longer than the other leg. What are the dimensions of the triangle. The answer is 8,15 and 17 - but I can't figure out how they did it. Can you show me?
Answer by richwmiller(17219)

a^2+b^2=c^2
a+9=c
b+2=c
that works out to
a=8
b=15
c=17
a=c-9
b=c-2
(c-9)^2+(c-2)^2=c^2
which yields
2c^2-22 c+85 = c^2
which gives 5 and 17 for c
solving for a and b
when c=5
a=c-9
b=c-2
a=5-9=-4
so 5 can't be solution since we can't have a negative number for a side of a triangle.
when c=17
a=17-9=8
b=17-2=15